OpenStudy (anonymous):

(2 & 4/5) (1 & 1/4) perform the indicated operation.

OpenStudy (anonymous):

what does & represent?

OpenStudy (anonymous):

they are mixed fractions. sorry.

OpenStudy (anonymous):

ahhh ok

OpenStudy (anonymous):

reduce to lowest terms.

OpenStudy (anonymous):

(2&4/5) = 2*5 + 4 = 14 (1&1/4) =1*4 +1 =5 so the answer is 14*5=70

OpenStudy (anonymous):

Ohhh so you would have to change them to mixed numbers meaning the numerator would be higher than the denominator. so for that you

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

Ahhh dang it lol Sorry

OpenStudy (anonymous):

35/8

OpenStudy (anonymous):

you mean that my answer is wrong?

OpenStudy (anonymous):

no, u just not done

OpenStudy (anonymous):

can you show the work. Of what remain please

OpenStudy (anonymous):

can I see it performed?

OpenStudy (anonymous):

\[2\frac{4}{5}= 2\times 5+4 = \frac{14}{5}\]

OpenStudy (anonymous):

Got you. Thanks

OpenStudy (anonymous):

So, 14/5 is the final answer? or 35/8?

OpenStudy (anonymous):

Can you teach me how to convert 14/5 into a mixed number?

OpenStudy (anonymous):

u wel com